OpenStudy (anonymous):

Prove that there is no rational number r such that r^3=2. I know that I can solve this by contradiction which uses the same argument that proves that there is no rational r for r^2=2 but my professor wants this solved two different ways so what's another way I can solve this?

OpenStudy (princeharryyy):

hope this helps @kdekle

OpenStudy (anonymous):

In all honesty it doesn't. I don't see how that helps me solve the problem.

OpenStudy (cwrw238):

i dont know if this makes sense but 1.25^3 = 1.953 1.26^3 = 2.000376 so the cube root of 2 lies between these 2 values can they possibly be a decimal number which ends in 1 - 9 with a cube which is a whole number? it seems that way - but i dont suppose this is a rigorous mathematical proof.

OpenStudy (anonymous):

You could use the rational roots test. If r were a rational number such that $$r^{3}=2,$$then r would be a rational root of the polynomial $$x^3-2=0.$$However, by the rational roots test, the only possible rational roots of this polynomial are 2 and -2, both of which aren't roots. Hence cube root of 2 can't be rational.

OpenStudy (anonymous):

Thank you that's awesome!